# A Survey on Mixture of Experts

### v1

@@ 5.1. Computation
Despite MoE is designed to scale model parameters efficiently without increasing computational demand, it encounters challenges pertaining to computational efficiency. One concern is the imbalance of computational load across distributed devices employing expert parallelism, which incurs significant synchronization overhead as the system awaits the processing completion of the most heavily loaded expert. Such issues are typically addressed through algorithmic strategies, such as optimized gating mechanisms and expert capacity adjustments, as discussed in the preceding section. Besides, solutions like SE-MoE (Shen et al., 2022), Tutel (Hwang et al., 2023a), FlexMoE (Nie et al., 2023) and SmartMoE (Zhai et al., 2023) have introduced dynamic expert placement strategies to distribute the workload as equally as possible among devices. Additionally, FasterMoE (He et al., 2022) has implemented a novel dynamic shadowed expert strategy, replicating experts on multiple devices to mitigate severe load imbalance. These model placement related strategies impact both computation and communication efficiency.
@@
Another concern is that MoE introduces additional computational overhead through operations including gate routing, input encode and output decode. Unlike expert computations, which mirror operations in dense models and benefit from extensive optimization on prevalent hardware such as GPUs, these MoE operations are characterized by redundant computation and memory movement, resulting in low efficiency on computing devices. Therefore, recent studies like DeepSpeed-MoE(Rajbhandari et al., 2022), FastMoE (He et al., 2021), HetuMoE (Nie et al., 2022) and Tutel (Hwang et al., 2023a) have focused on the development of tailored GPU kernels to enhance the efficiency of MoE operations.
@@
In contexts where multiple experts are deployed on a single GPU device, MegaBlocks (Gale et al., 2023) reformulates MoE computation in terms of block-sparse operations, developing specialized block-sparse GPU kernels that efficiently handle the dynamic workloads without dropping tokens. Zheng et al. (Zheng et al., 2023) propose PIT, a deep-learning compiler tailored for dynamic sparsity of MoE, which can find feasible PIT rules for all the operators within a model and generate optimized GPU kernels for them. PIT employs a novel tiling mechanism, utilizing the Permutation Invariant Transformation (PIT)—a mathematically proven property-to transform multiple sparsely located micro-tiles into a GPU-efficient dense tile without changing the computation results, thus achieving both high GPU utilization and low coverage waste. Despite these advancements, Tan et al. (Tan et al., 2024) highlight remaining optimization potential within current MoE frameworks such as MegaBlocks and PIT, which commence with an initial scatter-to-group data copy that increases memory footprint and requires a translation of the MoE problem into the sparse matrix format. Although this translation contributes minimally to computation overhead, it imposes limitations on the transparency and adaptability of extending MegaBlocks to modules beyond the FFN. To address these issues, Tan et al. (Tan et al., 2024) propose ScatterMoE, a MoE implementation designed to effectively minimize the memory footprint. ScatterMoE leverages ParallelLinear, a linear module capable of executing grouped matrix operations on scattered groups. This approach yields intermediate representations (e.g., the hidden states of an SMoE MLP) that are directly accessible as standard PyTorch tensors, allowing for easy extensions of MoE methods to other types of expert modules.
5.2. Communication
@@
To optimize the use of high intra-node bandwidth and low inter-node bandwidth, DeepSpeed-MoE (Rajbhandari et al., 2022) and HetuMoE (Nie et al., 2022) have introduced a hierarchical All-to-All communication strategy that enhances intra-node process and reduces inter-node data exchanges. Besides, FasterMoE (He et al., 2022), TA-MoE (Chen et al., 2022c) and SE-MoE (Shen et al., 2022) have introduced topology-aware routing strategies aimed at mitigating cross-node expert selection, thereby reducing inter-node communication burdens. Additionally, ExFlow (Yao et al., 2024) exploits expert affinity, anticipating expert allocation across layers to maximize the retention of token processing within local GPU confines. The strategic allocation of experts to minimize network traffic and leverage high-bandwidth connections is a prevalent approach in distributed MoE system (Rajbhandari et al., 2022; Singh et al., 2023; Tianwen Wei, 2024). And this is often integrated with the placement design of non-expert modules to optimize overall system performance.
@@
Given the concurrent feature of communication and computation, pipelining (Huang et al., 2019; Narayanan et al., 2019; Qi et al., 2023) is commonly employed to overlap their execution, thereby reducing the total time cost. This technique, which is integrated in systems such as Tutel (Hwang et al., 2023a), FasterMoE (He et al., 2022), and MPipeMoE (Zhang et al., 2024), orchestrates overlapping between All-to-All communication and expert computation. Notably, Lancet (Jiang et al., 2024b) underscores the inherent constraints of these pipelining methods, particularly the bounded duration for which expert computation and communication can overlap. To address this limitation, Lancet partitions non-MoE computations and integrates them into the pipeline during forward pass, and strategically schedules gradient weight computations to augment overlap in the backward pass. With the same objective of extending the overlap duration, ScMoE (Cai et al., 2024) restructures the MoE architecture to simultaneously process representations from preceding layers while engaging with current-layer representations. This decoupling of communication dependencies facilitates substantial, and in certain cases, complete overlapping between communication and computation. Snowflake Arctic (Team, 2024c) employs a similar design, utilizing a Dense-MoE hybrid transformer architecture to effectively overlap communication with computation.
5.3. Storage
@@
The ever-increasing parameters in MoE models exacerbate the constraints posed by memory capacity in compute devices, a challenge already pronounced in dense models. While expert parallelism offers a mitigation strategy through the distribution of experts across multiple devices, individual devices may still struggle to accommodate numerous experts, particularly in inference contexts where device capacity—such as that of edge devices (PCs, smartphones, IoTs)—is inherently more restricted.
@@
Considering the hierarchical storage pyramid, solutions like SE-MoE (Shen et al., 2022), Pre-gated MoE (Hwang et al., 2023b), and EdgeMoE (Yi et al., 2023) selectively retain only essential non-expert parameters and the active expert parameters within the GPU’s High-Bandwidth Memory (HBM), offloading inactive expert parameters to CPU memory or SSDs. These patterns incur additional overhead from data transfer across the storage hierarchy, thus they integrate expert selection forecasting and expert parameter prefetching techniques to overlap parameter access with computation.
@@
In addition, MPipeMoE (Zhang et al., 2024) introduces a strategy to reduce the memory overhead associated with activations and temporary buffers. This is achieved by sharing buffer for various partitions of tensors, while leveraging recomputation/communication and CPU offloading to recover the requisite activations in the backward pass.
6. Applications of Mixture of Experts Models
@@
In the current landscape dominated by Transformer-based large language models (LLMs), the mixture of experts (MoE) paradigm offers a compelling method to significantly expand model capacity while avoiding a corresponding surge in computational demands during training and inference phases. These models have been instrumental in enhancing the performance of LLMs across a spectrum of downstream tasks, with some applications achieving results that eclipse human performance (Fedus et al., 2022a; Jiang et al., 2024a; Dai et al., 2024). Rumors suggest that the formidable GPT-4 may employ an MoE architecture with an array of B experts, trained on diverse datasets and tasks, and utilizing a 16-iteration inference process 111https://x.com/soumithchintala/status/1671267150101721090. Given these, MoE has garnered widespread adoption across fields such as natural language processing, computer vision, recommender systems, and multimodal applications. The essence of these applications lies in leveraging conditional computation to significantly boost the number of model parameters, thereby augmenting model capacities with a fixed computational cost, or implementing dynamic expert selection through gating mechanisms for efficient multi-task learning. In the following, we will explore several representative applications of MoE in various domains, to provide a overall understanding of how MoE can be utilized to specific tasks.
@@
Natural Language Processing. The integration of MoE architectures with LLMs has unlocked extraordinary capabilities in a range of natural language understanding (NLU) and generation (NLG) tasks, including machine translation (Shazeer et al., 2017; Costa-jussà et al., 2022), open-domain question answering (Du et al., 2022; Artetxe et al., 2021), code generation (Jiang et al., 2024a; Dai et al., 2024; Team, 2024a; Tianwen Wei, 2024), and mathematical problem-solving (Dai et al., 2024; Jiang et al., 2024a; DeepSeek-AI, 2024; Team, 2024a). The methods of integrating MoE into LLMs have been thoroughly discussed and analyzed in the preceding algorithm design section 4 and system design section 5, and will not be reiterated in depth here.
@@
Computer Vision. The great success of sparsely-gated Mixture of Experts networks (MoE) in NLP has inspired their application in computer vision. For example, Riquelme et al. (Riquelme et al., 2021) introduced Vision MoE (V-MoE), which incorporates a sparsely activated mixture of MLPs into select ViT (Dosovitskiy et al., 2020) blocks. In image recognition tasks, V-MoE rivals the performance of state-of-the-art networks while requiring substantially less computational power during inference. This demonstrates the potential of MoE to discern distinct image semantics through specialized experts. Hwang et al. (Hwang et al., 2023a) develop Tutel, a scalable stack design and implementation for MoE with dynamic parallelism and pipelining, which they demonstrate with SwinV2-MoE, built upon Swin Transformer V2 (Liu et al., 2021). Moreover, Zhang et al. (Zhang et al., 2023) explore adversarial robustness in CNN-based MoE models, proposing a novel router-expert alternating adversarial training framework called ADVMOE. In most recent work, Chowdhury et al. (Chowdhury et al., 2023) introduce the concept of patch-level routing in MoE (pMoE) that segments each input image into patches (or tokens) and allocates patches () to each expert for processing through prioritized routing to enhance efficiency.
@@
Recommender System. Recommender systems are quintessential in various large-scale applications where they are required to balance and optimize multiple objectives simultaneously (Zheng and Wang, 2022). A prime example is in the domain of movie recommendations, where the aim is not only to suggest movie that align with users’ immediate preferences but also to ensure subsequent user satisfaction for the selected movies (Ma et al., 2018). The effectiveness of multi-task models hinges on the intricate interplay between task-specific goals and the relationships between tasks. Consequently, understanding the trade-offs inherent in these relationships is crucial. Mixture-of-experts (MoE) models with gating mechanisms have emerged as a popular paradigm for tackling the complexities of multi-task learning in recommender systems. Ma et al. (Ma et al., 2018) introduce the multi-gate mixture-of-experts (MMOE) approach, which capitalizes on the concept of shared expert submodels across all tasks, guided by a gating network tailored to each individual task. Addressing the “seesaw phenomenon” where the improvement of one task’s performance can detrimentally affect another is another challenge in multi-task learning. To counteract this, Tang et al. (Tang et al., 2020) propose the Progressive Layered Extraction (PLE) model for personalized recommendations. PLE distinctly segregates shared and task-specific components and employs a progressive routing mechanism to incrementally extract and refine the semantic knowledge, thereby enhancing the efficacy of joint representation learning and the routing of information across tasks. Recently, in the pursuit of capturing both the long-term and short-term user preferences that are particularly salient in sequential recommendation scenarios, a novel method named AdaMCT (Jiang et al., 2023b) has been proposed. AdaMCT utilizes layer-aware adaptive mixture units to dynamically blend CNN and Transformer experts, thereby tailoring the recommendations to individual user patterns.
@@
Multimodal Applications. Multimodal models are designed to process and integrate various data types within a single neural network framework (Ngiam et al., 2011). These models often simultaneously encompass two primary data modalities: images and text (Baltrušaitis et al., 2018; Uppal et al., 2022; Zhou et al., 2020). The Mixture of Experts (MoE) architecture has gained considerable traction as the foundation of multimodal models due to its capacity for expert layers to learn distinct modality partitioning (Mustafa et al., 2022). One notable implementation of this approach is the LIMoE model (Mustafa et al., 2022), a sparse mixture of expert models tailored for multimodal learning. LIMoE is trained on both images and text data, employing contrastive loss and an entropy-based regularization technique to address load balancing challenges inherent in MoE systems. Subsequently, Shen et al. (Shen et al., 2023b) and Lin et al. (Lin et al., 2024) have further investigated the potential of MoE for scaling vision-language models, offering valuable insights that contribute to the development of more efficient and effective multimodal learning systems. Furthermore, to address the specific issue of task conflicts in instruction tuning of Large Vision-Language Models (LVLMs), MoCLE (Gou et al., 2023) integrates MoE with LoRA (Hu et al., 2021) experts and a distinct universal expert to activate task-specific model parameters based on clusters of instructions. In parallel, to mitigate data conflicts, LLaVA-MoLE (Chen et al., 2024) deploys a set of LoRA experts, specifically for the MLP layer, combined with a top-1 gating mechanism to refine instruction tuning in Multimodal Large Language Models (MLLMs). While the MLLMs employing MoE architectures have demonstrated impressive performances, they generally involve a limited number of experts and modalities (Li et al., 2024a). To address this limitation, Li et al. (Li et al., 2024a) introduce the pioneering Uni-MoE, a unified MLLM with MoE architecture capable of managing an extensive range of modalities. They introduce a progressive training strategy to bolster expert collaboration and generalization across modalities, and they utilize LoRA (Hu et al., 2021), a lightweight fine-tuning methodology, to minimize computational demands.
7. Challenges & Opportunities

### v2

@@ 5.1 Computation
Despite MoE is designed to scale model parameters efficiently without increasing computational demand, it encounters challenges pertaining to computational efficiency. One concern is the imbalance of computational load across distributed devices employing expert parallelism, which incurs significant synchronization overhead as the system awaits the processing completion of the most heavily loaded expert. Such issues are typically addressed through algorithmic strategies, such as optimized gating mechanisms and expert capacity adjustments, as discussed in Section 4.1. Besides, solutions like SE-MoE [133], Tutel [132], FlexMoE [137] and SmartMoE [138] have introduced dynamic expert placement strategies to distribute the workload as equally as possible among devices. Additionally, FasterMoE [134] has implemented a novel dynamic shadowed expert strategy, replicating experts on multiple devices to mitigate severe load imbalance. These model placement related strategies impact both computation and communication efficiency.
@@
Another concern is that MoE introduces additional computational overhead through operations including gate routing, input encode, and output decode. Unlike expert computations, which mirror operations in dense models and benefit from extensive optimization on prevalent hardware such as GPUs, these MoE operations are characterized by redundant computation and memory movement, resulting in low efficiency on computing devices. Therefore, recent studies like DeepSpeed-MoE[66], FastMoE [131], HetuMoE [136] and Tutel [132] have focused on the development of tailored GPU kernels to enhance the efficiency of MoE operations.
@@
In contexts where multiple experts are deployed on a single GPU device, MegaBlocks [139] reformulates MoE computation in terms of block-sparse operations, developing specialized block-sparse GPU kernels that efficiently handle the dynamic workloads without dropping tokens. Zheng et al. [141] propose PIT, a deep-learning compiler tailored for dynamic sparsity of MoE, which can find feasible PIT rules for all the operators within a model and generate optimized GPU kernels for them. PIT employs a novel tiling mechanism, utilizing the Permutation Invariant Transformation (PIT)—–a mathematically proven property—to transform multiple sparsely located micro-tiles into a GPU-efficient dense tile without changing the computation results, thereby achieving both high GPU utilization and low coverage waste. Despite these advancements, Tan et al. [140] highlight remaining optimization potential within current MoE frameworks such as MegaBlocks and PIT, which commence with an initial scatter-to-group data copy that increases memory footprint and requires a translation of the MoE problem into the sparse matrix format. Although this translation contributes minimally to computation overhead, it imposes limitations on the transparency and adaptability of extending MegaBlocks to modules beyond the FFN. To address these issues, Tan et al. [140] propose ScatterMoE, a MoE implementation designed to effectively minimize the memory footprint. ScatterMoE leverages ParallelLinear, a linear module capable of executing grouped matrix operations on scattered groups. This approach yields intermediate representations (e.g., the hidden states of an SMoE MLP) that are directly accessible as standard PyTorch tensors, allowing for easy extensions of MoE methods to other types of expert modules.
5.2 Communication
@@
To optimize the use of high intra-node bandwidth and low inter-node bandwidth, DeepSpeed-MoE [66], HetuMoE [136] and ScheMoE [147] have introduced hierarchical All-to-All communication strategies that enhance intra-node process and reduce inter-node data exchanges. Besides, FasterMoE [134], TA-MoE [143] and SE-MoE [133] have introduced topology-aware routing strategies aimed at mitigating cross-node expert selection, thereby reducing inter-node communication burdens. Additionally, ExFlow [142] exploits expert affinity, anticipating expert allocation across layers to maximize the retention of token processing within local GPU confines. The strategic allocation of experts to minimize network traffic and leverage high-bandwidth connections is a prevalent approach in distributed MoE system [66, 135, 67]. Moreover, this is often integrated with the placement design of non-expert modules to optimize overall system performance.
@@
Given the concurrent feature of communication and computation, pipelining [205, 206, 207] is commonly employed to overlap their execution, thereby reducing the total time cost. This technique, which is integrated in systems such as Tutel [132], FasterMoE [134], PipeMoE [146] and MPipeMoE [144], orchestrates overlapping between All-to-All communication and expert computation. Notably, Lancet [145] underscores the inherent constraints of these pipelining methods, particularly the bounded duration for which expert computation and communication can overlap. To address this limitation, Lancet partitions non-MoE computations and integrates them into the pipeline during forward pass, and strategically schedules gradient weight computations to augment overlap in the backward pass. Punniyamurthy et al. [148] also emphasize the challenge posed by collective communications, which are often on the critical path, noting the difficulty of hiding their latency by overlapping kernel-granular communication and computation due to the absence of independent computation. Their solution involves fusing computation with dependent collective communication by leveraging GPU’s massive parallelism and GPU-initiated communication.
Aiming to break the inherent dependencies and thereby extend the overlap duration, ScMoE [110] restructures the MoE architecture to simultaneously process representations from preceding layers while engaging with current-layer representations. This decoupling of communication dependencies facilitates substantial, and in certain cases, complete overlapping between communication and computation. Snowflake Arctic [32] employs a similar design, utilizing a Dense-MoE hybrid transformer architecture to overlap communication with computation.
5.3 Storage
@@
The ever-increasing parameters in MoE models exacerbate the constraints posed by memory capacity in compute devices, a challenge already pronounced in dense models. While expert parallelism offers a mitigation strategy through the distribution of experts across multiple devices, individual devices may still struggle to accommodate numerous experts, particularly in inference contexts where device capacity—–such as that of edge devices (PCs, smartphones, IoTs)–—is inherently more restricted.
@@
Considering the hierarchical storage pyramid, solutions like SE-MoE [133], Pre-gated MoE [149], and EdgeMoE [150] selectively retain only essential non-expert parameters and the active expert parameters within the GPU’s High-Bandwidth Memory (HBM), offloading inactive expert parameters to CPU memory or SSDs. These patterns incur additional overhead from data transfer across the storage hierarchy, thus they integrate expert selection forecasting and expert parameter prefetching techniques to overlap parameter access with computation.
@@
In addition, MPipeMoE [144] introduces a strategy to reduce the memory overhead associated with activations and temporary buffers. This is achieved by sharing buffer for various partitions of tensors, while leveraging recomputation/communication and CPU offloading to recover the requisite activations in the backward pass.
6 Applications of Mixture of Experts
@@
In the current landscape dominated by Transformer-based large language models (LLMs), the mixture of experts (MoE) paradigm offers a compelling method to significantly expand model capacity while avoiding a corresponding surge in computational demands during training and inference phases. These models have been instrumental in enhancing the performance of LLMs across a spectrum of downstream tasks, with some applications achieving results that eclipse human performance [34, 29, 69]. Rumors suggest that the formidable GPT-4 may employ an MoE architecture with an array of B experts, trained on diverse datasets and tasks, and utilizing a 16-iteration inference process 111https://x.com/soumithchintala/status/1671267150101721090. Given these, MoE has garnered widespread adoption across fields such as natural language processing, computer vision, recommender systems, and multimodal applications. The essence of these applications lies in leveraging conditional computation to significantly boost the number of model parameters, thereby augmenting model capacities with a fixed computational cost, or implementing dynamic expert selection through gating mechanisms for efficient multi-task learning. In the following, we explore several representative applications of MoE in various domains to provide an overall understanding of how MoE can be utilized to specific tasks.
@@
Natural Language Processing. The integration of MoE architectures with LLMs has unlocked extraordinary capabilities in a range of natural language understanding (NLU) and generation (NLG) tasks, including machine translation [27, 78], open-domain question answering [35, 103], code generation [29, 69, 164, 67], and mathematical problem-solving [69, 29, 33, 164]. The methods of integrating MoE into LLMs have been thoroughly discussed and analyzed in Section 4 (algorithm design) and Section 5 (system design), and will not be reiterated in depth here. Beyond augmenting LLM capabilities, MoE has been instrumental in enhancing LLM’s safety while preserving its usability. A notable implementation is MoGU [151], which leverages dynamic routing to balance the contribution between usable LLM and safe LLM.
@@
Computer Vision. The great success of sparsely-gated Mixture of Experts networks (MoE) in NLP has inspired their application in computer vision. For example, Riquelme et al. [8] introduced Vision MoE (V-MoE), which incorporates a sparsely activated mixture of MLPs into selected ViT [211] blocks. In image recognition tasks, V-MoE rivals the performance of state-of-the-art networks while requiring substantially less computational power during inference. This demonstrates the potential of MoE to discern distinct image semantics through specialized experts. Hwang et al. [132] develop Tutel, a scalable system design and implementation for MoE with dynamic parallelism and pipelining, which they demonstrate with SwinV2-MoE, built upon Swin Transformer V2 [9]. Moreover, Zhang et al. [100] explore adversarial robustness in CNN-based MoE models, proposing a novel router-expert alternating adversarial training framework called ADVMOE. In most recent work, Chowdhury et al. [99] introduce the concept of patch-level routing in MoE (pMoE) that segments each input image into patches (or tokens) and allocates patches () to each expert for processing through prioritized routing to enhance efficiency.
@@
Recommender System. Recommender systems are quintessential in various large-scale applications where they are required to balance and optimize multiple objectives simultaneously [212]. A prime example is in the domain of movie recommendations, where the aim is not only to suggest movie that align with users’ immediate preferences but also to ensure subsequent user satisfaction for the selected movies [61]. The effectiveness of multi-task models hinges on the intricate interplay between task-specific goals and the relationships between tasks. Consequently, understanding the trade-offs inherent in these relationships is crucial. Mixture-of-experts (MoE) models with gating mechanisms have emerged as a popular paradigm for tackling the complexities of multi-task learning in recommender systems. Ma et al. [61] introduce the multi-gate mixture-of-experts (MMOE) approach, which capitalizes on the concept of shared expert submodels across all tasks, guided by a gating network tailored to each individual task. Addressing the “seesaw phenomenon” where the improvement of one task’s performance can detrimentally affect another is another challenge in multi-task learning. To counteract this, Tang et al. [152] propose the Progressive Layered Extraction (PLE) model for personalized recommendations. PLE distinctly segregates shared and task-specific components and employs a progressive routing mechanism to incrementally extract and refine the semantic knowledge, thereby enhancing the efficacy of joint representation learning and the routing of information across tasks. Recently, in the pursuit of capturing both the long-term and short-term user preferences that are particularly salient in sequential recommendation scenarios, a novel method named AdaMCT [153] has been proposed. AdaMCT utilizes layer-aware adaptive mixture units to dynamically blend CNN and Transformer experts, thereby tailoring the recommendations to individual user patterns. Zhang et al. [154] present MoE, an adaptive Multi-domain Multi-task Mixture-of-Experts framework designed for improving multi-domain multi-task recommendation. This framework incorporates a shared expert module, a domain expert module, and a task expert module to address the common information learning, domain-aspect user preferences, and task-aspect user preferences, respectively. Moreover, it employs a two-level fusion mechanism, powered by AutoML, ensuring precise control over feature extraction and fusion across diverse domains and tasks.
@@
Multimodal Applications. Multimodal models are designed to process and integrate various data types within a single neural network framework [213]. These models often simultaneously encompass two primary data modalities: images and text [214, 215, 216]. The Mixture of Experts (MoE) architecture has gained considerable traction as the foundation of multimodal models due to its capacity for expert layers to learn distinct modality partitioning [155]. One notable implementation of this approach is the LIMoE model [155], a sparse mixture of expert models tailored for multimodal learning. LIMoE is trained on both images and text data, employing contrastive loss and an entropy-based regularization technique to address load balancing challenges inherent in MoE systems. Subsequently, Shen et al. [156], Li et al. [157] and Lin et al. [106] have further investigated the potential of MoE for scaling vision-language models, offering valuable insights that contribute to the development of more efficient and effective multimodal learning systems. Furthermore, to address the specific issue of task conflicts in instruction tuning of Large Vision-Language Models (LVLMs), MoCLE [46] integrates MoE with LoRA [174] experts and a distinct universal expert to activate task-specific model parameters based on clusters of instructions. In parallel, to mitigate data conflicts, LLaVA-MoLE [115] deploys a set of LoRA experts, specifically for the MLP layer, combined with a top-1 gating mechanism to refine instruction tuning in Multimodal Large Language Models (MLLMs). While the MLLMs employing MoE architectures have demonstrated impressive performances, they generally involve a limited number of experts and modalities [158]. To address this limitation, Li et al. [158] introduce the pioneering Uni-MoE, a unified MLLM with MoE architecture capable of managing an extensive range of modalities. They introduce a progressive training strategy to bolster expert collaboration and generalization across modalities, and they utilize LoRA [174], a lightweight fine-tuning methodology, to minimize computational demands.
7 Challenges & Opportunities